OpenStudy (anonymous):

Nina's monthly budget is $2,250. Every month she makes the following payments: $175 for insurance $129 for utilities $283 for taxes Which is the highest house payment she can afford?Select one of the options below as your answer:A. House A - monthly payment: $1,580.89B. House B - monthly payment: $1,790.25C. House C - monthly payment: $1,989.23

OpenStudy (anonymous):

oops. A

OpenStudy (anonymous):

175+129+283= 587 2250 - 587 = 1663 so A

OpenStudy (anonymous):

2250 - (175 + 129 + 283) = 1663

OpenStudy (anonymous):

thank you :)